# XL-Instruct: Synthetic Data for Cross-Lingual Open-Ended Generation

### v1

@@ Abstract
Cross-lingual open-ended generation—i.e. generating responses in a desired language different from that of the user’s query—is an important yet understudied problem. We introduce XL-AlpacaEval, a new benchmark for evaluating cross-lingual generation capabilities in Large Language Models (LLMs), and propose XL-Instruct, a high-quality synthetic data generation method. Fine-tuning with just 8K XL-Instruct-generated instructions significantly improves model performance, increasing the win rate against GPT-4o-Mini from 7.4% to 21.5%, and improving on several fine-grained quality metrics. Additionally, models fine-tuned on XL-Instruct exhibit strong zero-shot transfer to both English-only and multilingual generation tasks. Given its consistent gains across the board, we strongly recommend incorporating XL-Instruct in the post-training pipeline of future multilingual LLMs. To facilitate further research, we will publicly and freely release the XL-Instruct and XL-AlpacaEval datasets, which constitute two of the few cross-lingual resources currently available in the literature.
@@ 1 Introduction
Cross-lingual generation is the task of understanding a query in a given source language and generating a response in a different target language. This task has assumed greater relevance in the recent development of Large Language Models (LLMs) with multilingual capabilities. Marchisio et al. (2024) noted its usefulness for a) companies that serve such LLMs across dozens of languages but optimizing a prompt for each input language is inefficient in practice, and b) when a user needs a generation in a language they do not speak. The conventional two-step ‘Reason then Translate’ approach to cross-lingual generation (Huang et al., 2023; Qin et al., 2023; Li et al., 2024b) can be problematic due to the noisy nature of machine translation (MT), which may lead to information loss or an unnatural-sounding response. It is also wasteful of inference time and cost, since the intermediary English response is thrown away once the desired cross-lingual output is obtained.
Despite its relevance, the adaptation of LLMs for cross-lingual generation is insufficiently studied. In the absence of high-quality instruction datasets and evaluation benchmarks, the cross-lingual abilities of LLMs remain emergent, yet inadequately assessed. In this work, we address the data deficiency for the cross-lingual generation task. We introduce XL-AlpacaEval, a cross-lingual evaluation dataset sourced from AlpacaEval (Li et al., 2023b), and observe poor off-the-shelf performance for most open multilingual LLMs. As a solution, we propose XL-Instruct, a synthetic data generation technique to create high-quality cross-lingual data at scale (illustrated in Figure 1) and show that fine-tuning with XL-Instruct significantly and consistently boosts cross-lingual performance across a range of base and instruction-tuned LLMs. Beyond cross-lingual capabilities, we demonstrate how this technique also shows strong zero-shot transfer performance in monolingual generation scenarios, both in English and in other languages. We will publicly release the XL-AlpacaEval benchmark and the XL-Instruct dataset, hoping to facilitate research in the cross-lingual LLM domain, which currently lacks sufficient resources for both evaluation and post-training.
We seek to answer the following Research Questions in this work:
  • RQ1: How good are SOTA multilingual LLMs on cross-lingual generation off-the-shelf? (§3)
  • RQ2: How well does XL-Instruct improve cross-lingual generation performance? (§3)
  • RQ3: How much does XL-Instruct transfer to zero-shot monolingual generation scenarios, both in English and non-English settings? (§6)
Figure 1: The XL-Instruct pipeline: 1) instruction generation from seed English data; 2) data refinement; 3) response translation into non-English; 4) data filtering, with more details in Section  4.
@@ 2 Related Work
Cross-Lingual Explorations in LLMs
Most of the current work on cross-lingual techniques in LLMs has focused primarily on prompting strategies . The popular approach is to leverage the extensive knowledge and superior reasoning capabilities of LLMs in high-resourced languages (like English) to improve performance in relatively lower-resourced ones (Qin et al., 2023; Huang et al., 2023; Singh et al., 2024; Wang et al., 2025). PLUG (Zhang et al., 2024) also shares this motivation, but fine-tunes an LLM to respond to a non-English question by first answering cross-lingually in English, followed by a translation to the target language. Extensions to this cross-lingual prompting paradigm have also emerged, such as X-InSTA (Tanwar et al., 2023) which uses a semantic encoder to select relevant cross-lingual examples, while SITR (Li et al., 2024b) employs self-reflection and iterative refinement to improve cross-lingual summarization. However, no prior study has approached cross-lingual open-ended generation as the primary training objective.
Synthetic Data Explorations
Previous studies on the creation of synthetic data for post-training LLMs has often been limited to English. Self-Instruct (Wang et al., 2023a) and Unnatural Instructions (Honovich et al., 2023) were among the earliest works to demonstrate how synthetic data pipelines could be used to instruction-tune an LLM. Some other efforts looked at the generation of diversified and high-quality data. Tülu 3 (Lambert et al., 2024), for instance, used persona-driven prompting to yield diverse synthetic data (Ge et al., 2024), while Llama 3 (Dubey et al., 2024) leveraged skill-specific experts as teacher models to generate data for coding, math, multilinguality, etc. To enable multilingual support, MT is often used to extend English resources to other languages (Muennighoff et al., 2023; Lai et al., 2023; Ranaldi & Pucci, 2023; Chen et al., 2024). Given the English resources are often model outputs themselves (eg. of chatGPT), training on translations of these can limit models’ exposure to diversity.
@@
There is a subset of data synthesis approaches called ‘reverse instruction’ methods which propose generating instructions from seed data, and then using the original seed data as responses to these instructions. Our work follows this trend of approaches. Initial works in this space Li et al. (2023a); Wang et al. (2023b) presented a two-step procedure which can be done iteratively: 1) fine-tuning a model to perform instruction generation, followed by 2) heuristic-based filtering to keep high-quality synthetic data. Additionally, Chen et al. (2023) also suggested “instruction wrapping” to refine response quality before fine-tuning the reverse instruction model. LongForm (Köksal et al., 2023) bypassed the fine-tuning stage in these works and leveraged a “teacher” LLM (InstructGPT) to generate such instructions off-the-shelf, yielding significant improvements in English text generation tasks. MURI (Köksal et al., 2024) and X-Instruction (Li et al., 2024a) extend LongForm to multilingual generation. The former back-translates to English, generates reverse instructions, and then forward-translates to low-resource languages. The latter bypasses back-translation to English and queries the teacher LLM in the low-resource language directly, potentially exposing the synthetic data to quality issues. The focus of these works is on improving ‘monolingual’ generation performance, ie. where query and response are in the same (non-English) language. Finally, Iyer et al. (2024a) and Iyer et al. (2024b) use similar strategies to create low-resource cross-lingual data for improving the translation performance of LLMs.
Unlike these previous works, our primary goal is to contribute data resources for cross-lingual open-ended generation, which includes a synthetic dataset where the instruction and response are in different languages, as well as a cross-lingual evaluation benchmark. [^1] To the best of our knowledge, we are the first work to propose a cross-lingual open-ended generation benchmark, and our synthetic training dataset is one of the few publicly available. Our experiments (see Table 9) show that it is of much higher quality than the closest prior work, X-Instruction (Li et al., 2024a). We intend to release the XL-Instruct dataset under permissive open source license.
@@ 3 XL-AlpacaEval: A Cross-Lingual Evaluation Benchmark
To benchmark cross-lingual open-ended generation, we create XL-AlpacaEval, which is adapted from the original AlpacaEval dataset (Li et al., 2023b). AlpacaEval contains 805 multi-domain prompts sampled from various test sets (Dubois et al., 2024), which includes OpenAssistant (Köpf et al., 2024), Koala (Geng et al., 2023), Vicuna (Chiang et al., 2023), Self-Instruct (Wang et al., 2023a) and Anthropic’s Helpfulness test set (Bai et al., 2022). Evaluation is carried out through the LLM-as-a-judge (Zheng et al., 2023) approach, wherein performance is measured by comparing the win rates of model outputs against reference (another model) outputs, as judged by an evaluator model.
To create XL-AlpacaEval, we first manually examine the AlpacaEval datasets and filter out prompts that are tailored towards eliciting responses in English. For example, questions about correcting grammar in an English sentence cannot be answered cross-lingually (refer Section A.1 for a detailed justification and list of prompts that were excluded). The filtered test set consists of 797 prompts. Next, we add cross-lingual generation instructions (such as “Answer in lang language”) to each of these prompts. We randomly sample these instructions from a list of templates (refer Section A.2 for a complete list) and create an evaluation set for 8 languages: 2 High-Resource EU including German (deu) and Portuguese (por), 2 Medium-Resource EU including Hungarian (hun) and Lithuanian (lit), 2 Low-Resource EU including Irish (gle) and Maltese (mlt), and 2 non-EU languages including Chinese (zho), and Hindi (hin). We also focus on evaluating the En-X direction in this work, as it might be the more common use case due to the prevalence of English text on the internet. It would be straightforward to extend this to other languages and pairs—one would simply need to run a script to append the cross-lingual templated instructions to our filtered test set.
@@
While the original implementation used GPT-4-Turbo as both reference and evaluator models, we use GPT-4o Mini for reference and GPT-4o as the judge, given the SOTA multilingual capabilities of the GPT-4o models. Our choice of using GPT-4o Mini as the reference model is motivated by two reasons: i) we experiment with ~7B-9B LLMs in this work, making the Mini model a suitable baseline; and ii) using different reference and evaluator models, with the more capable one assigned the role of the latter, should mitigate self-preference bias of models (Wataoka et al., 2024). Finally, GPT-4o has also been shown to obtain SOTA pairwise correlations with human preferences in multilingual chat scenarios (Gureja et al., 2024; Son et al., 2024) — making it suitable for our task.
Models
To evaluate the off-the-shelf cross-lingual capabilities of current multilingual LLMs, we benchmark several strong open-weight models in the 7B-9B parameters range: Aya Expanse 8B (Dang et al., 2024), Llama 3.1 8B Instruct (Dubey et al., 2024), Gemma 2 9B Instruct (Team et al., 2024), Qwen 2.5 7B Instruct (Yang et al., 2024), EuroLLM 9B Instruct (Martins et al., 2024), Aya 23 8B (Aryabumi et al., 2024) and Salamandra 7B Instruct (Gonzalez-Agirre et al., 2025). Inference of all models in this work is performed using the AlpacaEval repository (Li et al., 2023b), with the default decoding settings, i.e. temperature is 0.7, maximum tokens are set to 2048, and all models are loaded in bfloat16.
Results (Zero-Shot)
We show our benchmark scores in Table 1. Aya Expanse leads the table, achieving a  60% win rate against GPT-4o Mini for the four languages from X-AlpacaEval which it supports (por, deu, zho, hin). While it was trained on significant synthetic data itself using multilingual experts (Dang et al., 2024), it is not clear from the reportage whether its superiority is due to explicit cross-lingual tuning or implicit transfer. For the remaining languages, Llama 3.1 and Gemma 2 perform comparably with win rates ranging between 10% and 30%. We make two critical observations here. Firstly, except for Aya Expanse, most open LLMs trail significantly behind GPT-4o-Mini in cross-lingual generation, leaving significant room for improvement. Secondly, the performance varies widely based on the resourcefulness of the language. While Aya Expanse, Llama 3.1 and Gemma achieve win rates of 40% or higher for high-resource languages like por, deu and zho, performance drops to 20-30% for medium-resourced languages (hun, lit, hin) and even worse for lower-resourced languages like gle and mlt (10% or less). This suggests that we need a reliable pipeline to scalably create synthetic data across varied languages, without substantial quality drops for the lower-resourced ones, to ensure more consistent performances (see Table 6).
@@
Table 1: Win rates against GPT-4o (Mini) on XL-AlpacaEval, as judged by GPT-4o. Languages are denoted by their ISO 639-3 codes.
@@
Table 2: Performance changes on using Reason-then-Translate: scores represent differences against win rates from Table 1. Strong positive improvements are shaded.
Baseline: Reason-then-Translate
@@
Previous work had proposed prompting LLMs to reason first in a high-resource language such as English and then translating into the target language (Qin et al., 2023; Huang et al., 2023; Wang et al., 2025). We call this approach ‘reason-then-translate’ and report results in Table 2. The outcomes are mixed: stronger multilingual models like Aya Expanse, Llama, and Gemma suffer significant performance drops. Manual inspection reveals these 7B models occasionally produce empty outputs, likely due to difficulty consistently following complex instructions. We note that previous studies have only explored this strategy with larger ( 70B) models. In contrast, weaker LLMs like EuroLLM and Salamandra, fine-tuned mainly on English reasoning and MT data, can leverage this two-step approach to slightly improve their initially poor performance. Overall, our results indicate that improving cross-lingual generation performance is challenging and may not be easily resolved through prompting strategies alone.
4 The XL-Instruct Pipeline

### v2

# XL-Suite: Cross-Lingual Synthetic Training and Evaluation Data for Open-Ended Generation

@@ Abstract
Cross-lingual open-ended generation—responding in a language different from that of the query—is an important yet understudied problem. This work proposes XL-Instruct, a novel technique for generating high-quality synthetic data. We also introduce XL-AlpacaEval, a new benchmark for evaluating cross-lingual generation capabilities of large language models (LLMs). Our experiments show that fine-tuning with just 8K instructions generated using our XL-Instruct significantly improves model performance: increasing the win rate against GPT-4o-mini from 7.4% to 21.5% and improving on several fine-grained quality metrics. Moreover, base LLMs fine-tuned on XL-Instruct exhibit strong zero-shot improvements to same-language question answering, as shown on our machine-translated m-AlpacaEval. These consistent gains highlight the promising role of XL-Instruct in the post-training of multilingual LLMs. Finally, we publicly release XL-Suite, a collection of training and evaluation data to facilitate research in cross-lingual open-ended generation.
[…]
XL-Suite: Cross-Lingual Synthetic Training and Evaluation Data for Open-Ended Generation
Vivek Iyer1   Pinzhen Chen1,3   Ricardo Rei2,thanks: Work done while at Unbabel.   Alexandra Birch1 1University of Edinburgh  2Sword Health  3Queen’s University Belfast vivek.iyer@ed.ac.uk
@@ 1 Introduction
Cross-lingual generation is the task of understanding a query in a given source language and generating a response in a different target language. This task has assumed greater relevance in the recent era of large language models (LLMs) with multilingual capabilities. Marchisio et al. (2024) noted its usefulness for a) companies that serve such LLMs across dozens of languages, but optimizing a prompt for each input language is inefficient in practice, and b) when a user needs a generation in a language they do not speak. The conventional cascaded approaches to cross-lingual generation (Huang et al., 2023; Qin et al., 2023; Li et al., 2024b) could be problematic due to the noisy nature of machine translation, which leads to information loss or an unnatural-sounding response. It is also wasteful of inference time and cost, since the intermediary English response is thrown away once the desired cross-lingual output is obtained.
Figure 1: The XL-Instruct pipeline: 1) instruction generation from seed English data; 2) data refinement; 3) response translation into non-English; 4) data filtering, with more details in Section˜4.
The adaptation of LLMs to cross-lingual open-ended generation is relevant, given their versatile capabilities in both language conversion and question answering, but remains understudied. A primary reason is the absence of high-quality datasets and evaluation benchmarks. This work addresses the data deficiency for the cross-lingual generation task from both the modelling and evaluation perspectives. We first introduce XL-AlpacaEval, a cross-lingual evaluation benchmark built on AlpacaEval (Li et al., 2023b), and we observe poor off-the-shelf performance for most open-source multilingual LLMs. As a solution, we propose XL-Instruct, a synthetic data generation technique to create high-quality cross-lingual data at scale (illustrated in Figure 1) and show that fine-tuning with XL-Instruct significantly and consistently boosts cross-lingual performance across a range of base and instruction-tuned LLMs. Beyond cross-lingual capabilities, we also created a machine-translated benchmark for same-language generation, m-AlpacaEval, to demonstrate that our proposed data synthesis method achieves strong zero-shot transfer performance.
In this work, we seek to answer the following research questions through a comprehensive set of experiments:
  • RQ1: How good are off-the-shelf multilingual LLMs in cross-lingual generation? (§3)
  • RQ2: How does XL-Instruct improve cross-lingual capabilities of various LLMs? (§5.3)
  • RQ3: How does XL-Instruct fine-tuning impact standard multilingual same-language question answering? (§6)
Finally, to facilitate research in the cross-lingual LLM domain, which currently lacks sufficient resources for both evaluation and post-training, we publicly release XL-Suite [^1] https://huggingface.co/collections/viyer98/xl-suite-68ceb97cb1cc7e8499ffb971 – a comprehensive collection of cross-lingual training (XL-Instruct) and evaluation (XL-AlpacaEval and m-AlpacaEval) data.
@@ 2 Related Work
Cross-Lingual LLM Prompting
Most of the current research on cross-lingual generation in LLMs focuses on prompting strategies. The primary goal of generation here is to leverage the extensive knowledge and superior reasoning capabilities of LLMs in high-resourced languages (like English) to improve the final answer in lower-resourced ones (Qin et al., 2023; Huang et al., 2023; Singh et al., 2024; Wang et al., 2025). Similarly motivated, PLUG (Zhang et al., 2024) fine-tunes an LLM for this cross-lingual process: it first answers a non-English question by reasoning in English, then translates the response to the target language. Other extensions to this cross-lingual prompting paradigm have also emerged, such as X-InSTA (Tanwar et al., 2023), which uses a semantic encoder to select relevant cross-lingual examples, while SITR (Li et al., 2024b) employs self-reflection and iterative refinement to improve cross-lingual summarization. However, no prior study has approached cross-lingual open-ended generation as the primary training objective.
Data Synthesis
Previous studies on the creation of synthetic data for post-training LLMs have mostly been limited to monolingual scenarios, mostly in English. Self-Instruct (Wang et al., 2023a) and Unnatural Instructions (Honovich et al., 2023) were among the first to show how LLMs could be used to generate instructions from seed data. Later efforts have focused on generating diversified and skill-specific synthetic data. Tülu 3 (Lambert et al., 2024), for instance, used persona-driven prompting to yield diverse synthetic instructions (Ge et al., 2024), while Llama 3 (Dubey et al., 2024) leveraged skill-specific experts as teacher models to generate data for coding, math, multilinguality, etc. To enable multilingual support, machine translation is often used to extend English resources to other languages (Muennighoff et al., 2023; Lai et al., 2023; Ranaldi and Pucci, 2023; Chen et al., 2024). Given that English resources are often model outputs (e.g., of ChatGPT), training on translations of these can limit models’ exposure to diversity.
@@
A subset of data synthesis approaches relevant to our work is called “reverse instruction”, which generates instructions from seed data and then uses the original seed data as responses to these instructions, with back-translation being an early prominent example (Sennrich et al., 2016). Our work follows this trend of approaches applied to LLMs, where initial works (Li et al., 2023a; Wang et al., 2023b) presented a two-step procedure that can be done iteratively: 1) fine-tuning a model to perform instruction generation, followed by 2) heuristic-based filtering to keep high-quality synthetic data. Later, Chen et al. (2023) proposed “instruction wrapping” to refine response quality before fine-tuning the reverse instruction model. LongForm (Köksal et al., 2023) bypassed the fine-tuning step and leveraged a strong “teacher” LLM (InstructGPT) to generate such instructions directly, yielding significant improvements in English text generation tasks. MURI (Köksal et al., 2024) and X-Instruction (Li et al., 2024a) extend LongForm to multilingual generation. The former back-translates to English, generates reverse instructions, and then forward-translates to low-resource languages. The latter bypasses back-translation to English and queries the teacher LLM in the low-resource language directly, potentially exposing the synthetic data to quality issues. The focus of these works is on improving same-language generation performance. Finally, Iyer et al. (2024a) and Iyer et al. (2024b) use similar strategies to create low-resource cross-lingual data for boosting MT performance of LLMs.
Unlike these previous works, our primary goal is to contribute data resources for cross-lingual open-ended generation, which includes a synthetic dataset where the instruction and response are in different languages, as well as a cross-lingual evaluation benchmark. [^2] To the best of our knowledge, we are the first to propose a cross-lingual open-ended generation benchmark, and our synthetic training dataset is among the few publicly available. Our experiments (see Table 5) show that it is of much higher quality than the closest prior work, X-Instruction (Li et al., 2024a). We intend to release the XL-Instruct dataset under a permissive open source license.
@@ 3 XL-AlpacaEval: A Cross-Lingual Evaluation Benchmark
To evaluate cross-lingual open-ended generation, we create the XL-AlpacaEval benchmark, which is adapted from AlpacaEval v1 (Li et al., 2023b). AlpacaEval contains 805 multi-domain prompts sampled from various test sets (Dubois et al., 2024), including OpenAssistant (Köpf et al., 2024), Koala (Geng et al., 2023), Vicuna (Chiang et al., 2023), Self-Instruct (Wang et al., 2023a) and Anthropic’s Helpfulness test set (Bai et al., 2022). Evaluation is carried out through the LLM-as-a-judge approach (Zheng et al., 2023), where an evaluator LLM is used to estimate how often a model output would be preferred by humans over a baseline reference.
To create XL-AlpacaEval, we first manually examine the AlpacaEval dataset and filter out prompts that are tailored towards eliciting responses in English. For example, questions about correcting grammar in an English sentence cannot be answered cross-lingually (refer to Section˜A.1.1 for a detailed justification and a list of excluded prompts). The filtered test set consists of 797 prompts. Next, we add cross-lingual generation instructions (such as “Answer in {language}”) to prompts randomly sampled from a list of templates (in Section˜A.1.2) and create an evaluation set for eight languages, spanning resource availability, writing script, and geographical location: German (deu), Portuguese (por), Hungarian (hun), Lithuanian (lit), Irish (gle), Maltese (mlt), simplified Chinese (zho), and Hindi (hin). We focus on the En-X direction in this work, as generating in non-English is usually more challenging for LLMs that are usually English-centric. It should be straightforward to extend our benchmark to other languages and pairs—by appending the cross-lingual templated instructions to our filtered test set.
@@
While the original implementation used GPT-4-turbo as both reference and evaluator models, we use GPT-4o-mini for reference and GPT-4o as the judge, given GPT-4o’s strong multilingual capabilities. Our choice of using GPT-4o-mini as the reference model is motivated by two reasons: 1) we experiment with ~7–9B LLMs in this work, making the GPT-4o-mini model a suitable baseline; and 2) using different reference and judge models, with the more capable one as the judge, should mitigate self-preference bias of models (Wataoka et al., 2024). Finally, GPT-4o has also been shown to obtain state-of-the-art pairwise correlations with human ratings in multilingual chat scenarios (Gureja et al., 2024; Son et al., 2024).
@@
Table 1: Zero-shot win rates against GPT-4o-mini on XL-AlpacaEval as judged by GPT-4o.
@@
Table 2: Performance change over zero-shot when using Reason-then-Translate: scores represent differences against win rates from Table˜1. Strong positive improvements are shaded.
Models
To evaluate off-the-shelf cross-lingual capabilities of existing multilingual LLMs, we benchmark several strong open-weight models in the ~7–9B parameter range: Aya Expanse 8B (Dang et al., 2024), Llama 3.1 8B Instruct (Dubey et al., 2024), Gemma 2 9B Instruct (Team et al., 2024), Qwen 2.5 7B Instruct (Yang et al., 2024), EuroLLM 9B Instruct (Martins et al., 2024), Aya 23 8B (Aryabumi et al., 2024), and Salamandra 7B Instruct (Gonzalez-Agirre et al., 2025). Inference is performed using the AlpacaEval repository (Li et al., 2023b), with the default decoding settings: temperature 0.7, maximum tokens 2048, and models loaded in bfloat16.
Zero-Shot Results
We show our benchmark scores in Table˜1. Aya Expanse leads the table, achieving a  60% win rate against GPT-4o-mini for the four languages it supports (por, deu, zho, hin). While it was trained on significant synthetic data using multilingual experts (Dang et al., 2024), it remains unclear whether its superiority stems from explicit cross-lingual tuning or implicit transfer. For other languages, Llama 3.1 and Gemma 2 yield comparable win rates ranging between 10% and 30%. We make two critical observations here. Firstly, except for Aya Expanse, most open LLMs trail significantly behind GPT-4o-mini in cross-lingual generation, leaving much room for improvement. Secondly, the performance strongly correlates with the resourcefulness of the language. While Aya Expanse, Llama 3.1, and Gemma achieve win rates of 40% or higher for high-resource languages like por, deu, and zho, performance drops to 20-30% for medium-resourced languages (hun, lit, hin) and 10% or less for lower-resourced languages like gle and mlt. This underscores the need for scalable pipelines for creating high-quality synthetic data for lower-resourced languages, in order to achieve more consistent model performance (see Table 9).
Reason-then-Translate Results
@@
Previous works have proposed prompting LLMs to reason first in a high-resource language (e.g., English) and then translating into the target language (Qin et al., 2023; Huang et al., 2023; Wang et al., 2025). We call this approach “reason-then-translate” and report results in Table 2. The outcomes are mixed: stronger multilingual models like Aya Expanse, Llama, and Gemma suffer significant performance drops. Manual inspection reveals these 7B models occasionally produce empty outputs, likely due to difficulty in following complex multi-step instructions—this aligns with prior findings which report successful results from only larger models (Hu et al., 2025). In contrast, weaker LLMs like EuroLLM and Salamandra, fine-tuned on English reasoning and MT data, can leverage this two-step approach to yield some gains over their poor initial scores. Overall, these results show that inducing cross-lingual capabilities in standard multilingual LLMs may not be resolved through prompting strategies alone.
4 The XL-Instruct Data Synthesis Pipeline